three years ago zery was three times as old as rick three years hence she will be twice as old.find their present age
Answer by lwsshak3(11628)

three years ago zery was three times as old as rick three years hence she will be twice as old.find their present age
***
3 yrs ago:
x=rick's age
3x=zery's age
..
at present:
x+3=rick's age
3x+3=zery's age
..
3 yrs hence
x+6=rick's age
3x+6=zery's age
3x+6=2(x+6)
3x+6=2x+12
x=6
present ages:
x+3=9
3x+3=21
rick's present age=9
zery's present age=21